OpenStudy (anonymous):

y'=ty(3-y) Describe how solutions appear to behave as t increases and how their behavior depends on the initial value y0 ("0" is a subscript) when t=0. How do I get the solutions in order to plot them?

OpenStudy (anonymous):

y'=dy/dt so you can simplify above equation as \[\frac{ dy }{ y(3-y)}=tdt\] now integrate this by using partial fraction \[\frac{ 1 }{ 3} [\frac{ 1 }{ y }+ \frac{ 1 }{ 3-y }]dy= tdt\] i think you can integrate it now

OpenStudy (anonymous):

I got that far. After that point this is what I have: Integrating gets me to: \[\frac{ 1 }{ 3 }\left( \ln \frac{ y }{3-y} \right)=1/2t ^{2}+C\] then simplifying and trying to solve for C, I get stuck at: \[\frac{ y }{ 3-y }=Ce ^{(3/2)t ^{2}}\] Not sure how to proceed from here to get C and then the solutions based on what the question is asking for.